OpenStudy (anonymous):

simplify completely 18st^4/52s^3t*16s^3/9s^2t

jhonyy9 (jhonyy9):

this is a product of two fractions ?

jhonyy9 (jhonyy9):

on the first step simplifie the first fraction try it and see what will get than the second one

OpenStudy (solomonzelman):

Putting up the question clearly. \[\huge\color{blue}{ \frac{18st^4}{52s^3t} } \times \huge\color{blue}{ \frac{16s^3}{9s^2t } } \]

OpenStudy (anonymous):

4t^2/13s

jhonyy9 (jhonyy9):

this is the result what you have got ?

OpenStudy (anonymous):

yes did i do it wrong? i can try again

jhonyy9 (jhonyy9):

make the calcules step by step

OpenStudy (anonymous):

can you write the steps out?

jhonyy9 (jhonyy9):

yes i can solve it but i like seeing how you have got this result

OpenStudy (anonymous):

oh no i dont want you to solve it i want u to tell me what to do so i can get the write answer

OpenStudy (anonymous):

oh i found what i did wrong the answer is 8t^2/13s

jhonyy9 (jhonyy9):

18st^4/52s^3t*16s^3/9s^2t 18 divide 52 simplifie by 2 see what will get and continue it with s simplifie by s on exponent 3 so what will get and t on exponent 4 simplifie it by t and see what wil get than after these continue these with second fraction

jhonyy9 (jhonyy9):

so you dont like rewriting these calculi how you have got this result ? but this answer is right

OpenStudy (anonymous):

thnk you. its just that it takes so much time

jhonyy9 (jhonyy9):

yw my pleasure